The acceleration of gravity at the surface of Mars is 3.7 m/s2. If a stone is thrown upward in Mars reaches a height of .01 km. Find its a) initial speed; b) the total time of flight; and c) its location after 1.5 sec.



1
Expert's answer
2021-09-28T07:30:11-0400

We can input the equation in: v = 0

It is informed to think that the speed at maximum height is 0; the stone goes up at a certain speed, stops, and then falls down) g = -3.7

(the acceleration is given in the problem) Δy = 19 (the height is given in the problem)

v0 (the initial velocity) v^2 = v0^2 + 2gΔy 0^2 = v0^2 + 2(-3.7)(19) 0 = v0^2 + (-140.6) 140.6 = v0^2